Exercise 2: Calculate consumer surplus for the demand curve P=1-Q and the supply curve P=1+Q 


1
Expert's answer
2021-07-30T03:48:01-0400

Given :

demand curve : "P = 1 - Q"

Supply curve : "P = 1 + Q"

Equilibrium is attained where demand equals the supply ie,

"1 - Q = 1 + Q"

"2Q = 0"

"Q = 0"

"P = 1"

Consumer Surplus = "\frac {1} {2} × Quantity ×( Consumer willingness to pay - Price actually paid)"


Consumer Surplus = "\frac {1} {2} × 0"


Consumer Surplus = 0


Thus, Consumer surplus is 0.